# AI-driven visual monitoring of industrial assembly tasks

### v1

@@ 1 Introduction
In this work, we introduce ViMAT, a novel AI-driven system for real-time Visual Monitoring of industrial Assembly Tasks that operates without such constraints. We formalize the problem as a state-transition system where states are specific configurations of assembly components and transitions between them occur depending on the actions performed by the human operator. ViMAT combines a perception module, which extracts visual observations from multi-view video streams, with a probabilistic reasoning module that infers the most likely action being performed based on the observed assembly state and prior task knowledge (Fig. 1). The perception module leverages a deep learning-based object detector trained exclusively on synthetic data generated from digital twins that replicate the physical assembly components and eliminate the need for manual annotation of real-world data. The effectiveness of the resulting detector eliminates the need for rigid workspace organization and visual markers, while its robustness to lighting variations enables accurate detection in real-world industrial settings. To mitigate occlusions, we capture video footage from multiple viewpoints, perform object detection independently for each viewpoint, and then fuse the resulting predictions to produce spatially coherent observations. The probabilistic reasoner uses these observations to predict the full sequence of actions leading to the current configuration and demonstrates robustness to ambiguous actions.
@@ 2 Related works
Visual monitoring of human-operated tasks aims to identify the actions performed by individuals from video recordings. In industrial environments, this capability is essential for ensuring adherence to standard operating procedures, promoting safe machinery usage, and safeguarding worker safety. Existing approaches typically interpret observed behavior using either rule-based reasoning or probabilistic models. In this work, we focus on the latter and address real-world scenarios where the partial observability caused by inherent occlusions poses a significant challenge. Most competitors [27, 31, 6, 32] rely on the interpretation of human actions, which can be ambiguous, deviate from standard procedures, or vary significantly across operators. In contrast, our approach centers on detecting assembly components and then infers the underlying action by identifying the predefined assembly instruction that is most likely to have resulted in the observed configuration.
Digital twins are virtual replicas of physical objects that mirror their real-world counterparts either in appearance alone or also in functionality [21, 24]. They are essential for enabling the analysis and simulation of complex systems or processes, particularly when direct inspection is impractical, costly, or hazardous. As a result, digital twins are playing an increasingly important role in manufacturing [8, 30], robotics [14, 26], and autonomous driving [19, 34]. More recently, they have become valuable tools for training, testing, and monitoring computer vision systems. In particular, they have been used in combination with rendering engines [1, 12] to generate synthetic datasets for training object detectors [18], pose estimators [22], and segmentation models [13]. A key advantage of using digital twins in this context is the automatic generation of accurate annotations, reducing the burden of manually labeling real-world data. In this work, we create digital twins of the assembly components, replicating both their shape and appearance. We extract precise 3D geometry from synthetic CAD models and approximate visual appearance using synthetic materials [2, 4] that simulate photometric properties (e.g., color, texture) and physical characteristics (e.g., reflectivity, signs of wear).
@@ 3 Method
Assembly tasks are formally defined by three components: objects, configurations, and actions. Objects include physical elements, which are manipulated during the assembly process, as well as trays, which are designated regions of the workspace associated with specific functions, such as storing elements before use, collecting unmounted elements, or holding the required tools. Note that we do not require trays to be physical containers. Configurations describe the relationships among objects and are defined by a set of Boolean predicates over one or more objects that hold true in a specific assembly state. Actions represent admissible interactions with objects that cause transitions between configurations by modifying the associated predicates. Every action is associated with a set of preconditions, i.e. predicates that must be satisfied for the action to be applicable. We model an assembly task as a state-transition system, where states correspond to configurations and transitions to actions [11, 16]. In this framework, a planner can identify the correct executions of an assembly task by finding ordered sequences of states transitioning from the input initial configuration $C_{\text{init}}$ to the target final configuration $C_{\text{final}}$ through valid actions. This system can be represented as a directed graph, where nodes correspond to configurations (with $C_{\text{init}}$ as root and $C_{\text{final}}$ as the sole leaf), edges represent actions, and paths connecting $C_{\text{init}}$ to $C_{\text{final}}$ define valid plans. Note that multiple such paths may exist, each corresponding to a different valid plan (see Fig. 3).
@@ 3.2 Perception
Object detection. We train an object detector $\Phi_{\Theta}$ on a synthetic dataset generated using a digital twin of the task-specific assembly environment (see Section 3.1). At inference time, $\Phi_{\Theta}$ performs object detection frame-wise, producing 2D bounding boxes and class predictions.
Multi-view fusion. To improve the robustness and accuracy of $\Phi_{\Theta}$, we leverage multiple cameras through a multi-view fusion strategy. The process begins by selecting, for each viewpoint, the 2D detections that fall within each predefined tray region. This results in a set of per-tray, per-view detections. These 2D detections are then lifted into 3D point clouds using the known camera intrinsics and extrinsics, by applying the pinhole camera model. To associate detections across different viewpoints, we match the 3D point clouds using a pairwise Point Cloud Intersection-over-Union (IoU) metric. This is based on a radius-based intersection strategy. Formally, given two point clouds $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, and a radius threshold $r$, we define the intersection set as: $\text{Intersection}(\mathcal{P}_{1},\mathcal{P}_{2},r)=\left\{p\in\mathcal{P}_{2}\;\middle|\;\exists q\in\mathcal{P}_{1}\text{ such that }\|p-q\|<r\right\}$. The corresponding IoU value is computed using the usual formula. This method allows us to robustly identify overlapping detections across views, even in the presence of minor inaccuracies. Once the pairwise IoUs are computed and matching detections are identified, the results are aggregated across all views to obtain a consolidated observation. This observation consists of the elements present in the workspace, each associated with a confidence score. The multi-view fusion helps mitigate occlusions and provides a more reliable perception of the scene.
Temporal smoothing. To ensure consistency over time and reduce noise in the predictions, we apply a temporal smoothing strategy to the observation vectors. At each time step, the current observation vector is combined with the previously smoothed one using a weighted average. The weighting is asymmetric: if the current score of an element is higher than the previous one, it is given more weight to allow rapid adaptation to improvements; otherwise, the previous score is favored to dampen abrupt drops.
@@ 4.1 Experimental setting
We conduct experiments in two distinct scenarios: a controlled LEGO scenario and an industrial scenario. In both scenarios, we use three RGB-D sensors for perception, and the communications between system modules are handled via sockets. We train as object detector $\Phi_{\Theta}$, the pretrained YOLOv8-X model (68.2M parameters) and finetune it for 100 epochs. We employ a cosine decay schedule, and the SGD optimizer. Input images are resized to 480×480 pixels with standard data augmentation techniques (random scaling, rotation, and HSV variations).
@@
4.2 Quantitative results
We present quantitative evaluations of ViMAT, focusing on object detection performance in both the LEGO and industrial scenarios. Tables 2 and 2 detail the average mAP results for each setting, highlighting the impact of data augmentation (Aug.) and background images (BG) during training.
@@ 4.2 Quantitative results
Industrial scenario. Tab. 2 shows the mAP results across the three cameras and their average. In this more challenging scenario, our method achieves a peak performance of 58.4%, reflecting the increased difficulty in detecting similar objects. The models trained using only real images perform worse than in the LEGO scenario, likely due to the limited training data (only 8 images). Tab. 2 includes ablation studies, such as variations using synthetic and real data (S+R), and SAM-6D [23] performance. Combining synthetic and real data for training (S+R) yields improved results, highlighting the benefit of including real data when available. SAM-6D, a model-based zero-shot method, achieves lower performance, likely due to its poor calibration.
4.3 Qualitative results
Figures 3 and 4 show the state probabilities over time for different videos. Each graph compares the ground-truth states (pink) with ViMAT’s predictions (blue or green) and with predictions obtained using ground-truth object detections as observation vectors (azure). Precision and recall are computed by matching predicted and ground-truth active states at each time step.
LEGO scenario. At the bottom of Fig. 3, the complete assembly path is shown, illustrating the system’s ability to track the progression from the initial (left) to the final configuration (right). ViMAT closely follows the ground-truth sequence of states, demonstrating strong temporal alignment. Notably, predictions from the trained object detector closely match those derived from oracle detections, indicating that the model, trained solely on synthetic data, generalizes well to the real-world and provides reliable input for downstream reasoning.
@@ 5 Conclusions
Acknowledgments. This work was supported by the PAT Legge 6 project NextMag. This work has been developed within a collaboration between FBK and Meccanica del Sarca s.p.a. and funded by the PAT Legge 6 project NextMag.

### v2

@@ 1 Introduction
In this work, we introduce ViMAT, a novel AI-driven system for real-time Visual Monitoring of industrial Assembly Tasks that operates without such constraints. We formalize the problem as a state-transition system, where states are specific configurations of assembly components, and transitions between them occur depending on the actions performed by the human operator. ViMAT combines a perception module, which extracts visual observations from multi-view video streams, with a probabilistic reasoning module that infers the most likely action being performed based on the observed assembly state and prior task knowledge (Fig. 1). The perception module leverages a deep learning-based object detector trained exclusively on synthetic data generated from digital twins that replicate the physical assembly components and eliminate the need for manual annotation of real-world data. The effectiveness of the resulting detector eliminates the need for rigid workspace organization and visual markers, while its robustness to lighting variations enables accurate detection in real-world industrial settings. To mitigate occlusions, we capture video footage from multiple viewpoints, perform object detection independently for each viewpoint, and then fuse the resulting predictions to produce spatially coherent observations. The probabilistic reasoner uses these observations to predict the full sequence of actions leading to the current configuration and demonstrates robustness to ambiguous actions.
@@ 2 Related works
Visual monitoring of human-operated tasks aims to identify the actions performed by individuals from video recordings. In industrial environments, this capability is essential for ensuring adherence to standard operating procedures, promoting safe machinery usage, and safeguarding worker safety. Existing approaches typically interpret observed behaviour using either rule-based reasoning or probabilistic models. In this work, we focus on the latter and address real-world scenarios where the partial observability caused by inherent occlusions poses a significant challenge. Most competitors [27, 31, 6, 32] rely on the interpretation of human actions, which can be ambiguous, deviate from standard procedures, or vary significantly across operators. In contrast, our approach centers on detecting assembly components and then infers the underlying action by identifying the predefined assembly instruction that is most likely to have resulted in the observed configuration.
Digital twins are virtual replicas of physical objects that mirror their real-world counterparts either in appearance alone or also in functionality [21, 24]. They are essential for enabling the analysis and simulation of complex systems or processes, particularly when direct inspection is impractical, costly, or hazardous. As a result, digital twins are playing an increasingly important role in manufacturing [8, 30], robotics [14, 26], and autonomous driving [19, 34]. More recently, they have become valuable tools for training, testing, and monitoring computer vision systems. In particular, they have been used in combination with rendering engines [1, 12] to generate synthetic datasets for training object detectors [18], pose estimators [22], and segmentation models [13]. A key advantage of using digital twins in this context is the automatic generation of accurate annotations, reducing the burden of manually labelling real-world data. In this work, we create digital twins of the assembly components, replicating both their shape and appearance. We extract precise 3D geometry from synthetic CAD models and approximate visual appearance using synthetic materials [2, 4] that simulate photometric properties (e.g., colour, texture) and physical characteristics (e.g., reflectivity, signs of wear).
@@ 3 Method
Assembly tasks are formally defined by three components: objects, configurations, and actions. Objects include physical elements, which are manipulated during the assembly process, as well as trays, which are designated regions of the workspace associated with specific functions, such as storing elements before use, collecting unmounted elements, or holding the required tools. Note that we do not require trays to be physical containers. Configurations describe the relationships among objects and are defined by a set of Boolean predicates over one or more objects that hold true in a specific assembly state. Actions represent admissible interactions with objects that cause transitions between configurations by modifying the associated predicates. Every action is associated with a set of preconditions, i.e. predicates that must be satisfied for the action to be applicable. We model an assembly task as a state-transition system, where states correspond to configurations and transitions to actions [11, 16]. In this framework, a planner can identify the correct executions of an assembly task by finding ordered sequences of states transitioning from the input initial configuration $C_{\text{init}}$ to the target final configuration $C_{\text{final}}$ through valid actions. This system can be represented as a directed graph, where nodes correspond to configurations (with $C_{\text{init}}$ as root and $C_{\text{final}}$ as the sole leaf), edges represent actions, and paths connecting $C_{\text{init}}$ to $C_{\text{final}}$ define valid plans. Note that multiple such paths may exist, each corresponding to a different valid plan (see Fig. 3, bottom).
@@ 3.2 Perception
Object detection. We train an object detector $\Phi_{\Theta}$ on a synthetic dataset generated using a digital twin of the task-specific assembly environment (see Section 3.1). At inference time, $\Phi_{\Theta}$ performs object detection frame-wise, producing 2D bounding boxes, class predictions, and confidence scores.
Multi-view fusion. To improve the robustness and accuracy of $\Phi_{\Theta}$, we leverage multiple cameras through a multi-view fusion strategy. The process begins by selecting, for each viewpoint, the 2D detections that fall within each predefined tray region. This results in a set of per-tray, per-view detections with associated confidence scores. Every pixel inside the 2D detection is back-projected using depth, intrinsics, and the pinhole model, then transformed to the global frame via extrinsics, yielding a 3D point cloud per detection per view. To associate detections across viewpoints, we match these point clouds using a pairwise Point Cloud Intersection-over-Union (IoU) metric. This is based on a radius-based intersection strategy. Formally, given two point clouds $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, and a radius threshold $r$, we define the intersection set as: $\text{Intersection}(\mathcal{P}_{1},\mathcal{P}_{2},r)=\left\{p\in\mathcal{P}_{2}\;\middle|\;\exists q\in\mathcal{P}_{1}\text{ such that }\|p-q\|<r\right\}$. The corresponding IoU value is computed using the usual formula. This method allows us to robustly identify overlapping detections across views, even in the presence of minor inaccuracies. Once the pairwise IoUs are computed and matching detections are identified, the results are aggregated across all views to obtain a consolidated observation, i.e., a vector representing the presence and confidence of each element in the workspace. The multi-view fusion helps mitigate occlusions and provides a more reliable perception of the scene.
Temporal smoothing. To ensure consistency over time and reduce noise in the predictions, we apply a temporal smoothing strategy to the consolidated observation vectors. At each time step, the current observation vector is combined with the previously smoothed one using a weighted average. The weighting is asymmetric: if the current score of an element is higher than the previous one, it is given more weight to allow rapid adaptation to improvements; otherwise, the previous score is favored to dampen abrupt drops.
@@ 4.1 Experimental setting
We conduct experiments in two distinct scenarios: a controlled LEGO scenario and an industrial scenario. In both scenarios, we use three RGB-D sensors for perception, and the communications between system modules are handled via sockets. We train as object detector $\Phi_{\Theta}$, the pretrained YOLOv8-X model (68.2M parameters) and finetune it for 100 epochs. We employ a cosine decay schedule, and the SGD optimizer. Input images are resized to 480$\times$480 pixels with standard data augmentation techniques (random scaling, rotation, and HSV variations).
@@
4.2 Results
Quantitative Results. We present quantitative evaluations of ViMAT’s object detection performance in both LEGO and industrial scenarios. Tables 2 and 2 detail the average mAP results for each setting, highlighting the impact of data augmentation (Aug.) and background images (BG) during training.
@@ 4.2 Results
Industrial scenario. Tab. 2 shows the mAP results across the three cameras and their average. In this more challenging scenario, our method achieves a peak performance of 58.4%, reflecting the increased difficulty in detecting similar objects. The models trained using only real images perform worse than in the LEGO scenario, likely due to the limited data (only 8 images). Tab. 2 includes ablation studies, such as variations using synthetic and real data (S+R), and SAM-6D [23] performance. Combining synthetic and real data for training (S+R) yields improved results, highlighting the benefit of including real data when available. SAM-6D, a model-based zero-shot method, achieves lower performance, likely due to its poor calibration.
Qualitative Results. Figures 3 and 4 show the state probabilities over time for different videos. Each graph compares the ground-truth states (pink) with ViMAT’s predictions (blue or green) and with predictions using ground-truth object detections as observation vectors (azure). Precision and recall are computed by matching predicted and ground-truth active states at each time step.
LEGO scenario. At the bottom of Fig. 3, the complete assembly path is shown, illustrating the system’s ability to track the progression from the initial (left) to the final configuration (right). ViMAT closely follows the ground-truth sequence, demonstrating strong temporal alignment. Notably, predictions from the trained object detector closely match those derived from oracle detections, indicating that the model, trained solely on synthetic data, generalizes well to the real-world and provides reliable input for downstream reasoning.
@@
Failure cases and limitations. A key limitation is the perception module, where missed or incorrect detections, especially for similar objects, impact downstream reasoning. The system also struggles to distinguish between actions with subtle differences.
@@ 5 Conclusions
Acknowledgments. This work has been partially funded by the Provincia Autonoma di Trento (Italy) under L.P. 6/99, as part of the NEXTMAG project.